OpenStudy (anonymous):

Differentiate Cos^-2(x)

OpenStudy (anonymous):

Use the Chain Rule for inner function \(g(x)=\cos x\) and outer function \(f(x)=\frac{1}{x^2}\). \[f(g(x)) = \left(\cos x\right)^{-2}\]

OpenStudy (anonymous):

i thought u use the inverse rule

OpenStudy (anonymous):

I'm not familiar with any inverse rule?

OpenStudy (anonymous):

Presumably you are differentiating with respect to x? I hate to be picky, but you should get use to specifying that. Later on (higher math courses and/or more involved problems) it won't be so obvious. y =( cosx)^-2 dy/dx = -2[(cosx)^-3] (cosx)' dy/dx = -2[cosx)^-3] (-sinx) dy/dx = 2 [(cosx)^-3] sinx dy/dx = 2 sinx (secx)^3